OpenStudy (aliza29faye):

A sky diver accelerates until the drag force (friction) of air causes him to reach what is called terminal velocity. At that point he falls at a constant rate of around 55 meters per second (m/s). The weight of the skydiver on the ground is 800 N. What is the drag force at terminal velocity?

OpenStudy (anonymous):

From the inertia's law, a body moves with constant velocity if and only if the sum of the forces acting on it is 0. You know the magnitude of one force, and you know that only another force is acting (friction), so the magnitude of friction is the solution of 800 + x = 0.

OpenStudy (anonymous):

let the drag force on the diver is Fd Now , when the diver achieves terminal velocity , as the velocity is now constant , acceleration of diver is zero Using second law of motion net force = m*a Fd - m*g = 0 Fd - 800 = 0 Fd = 800 N the drag force acting on diver is 800 N in upwards direction